OpenStudy (anonymous):

For this set of simultaneous equations: (m-3)x+8y=10 and 5x+(m+3)y=11, find the values of m for which the solutions have: no solutions, unique solutions and infinite solutions.

OpenStudy (anonymous):

I have found that the values for m are 7 and -7 by substitution and then equating co-efficients.

OpenStudy (anonymous):

* no solutions, unique solutions and infinite solutions. The key is first understanding what those terms mean.

OpenStudy (anonymous):

I do. No solutions means that the graphs don't intersect; unique solutions means that there is a set amount of solutions and infinite solutions means that there are a LOT of solutions.

OpenStudy (anonymous):

I got the solutions for m and I don't know what do with them.

OpenStudy (anonymous):

If you try to solve in terms of m, you will get \[ \left( \begin{array}{cc} x=\frac{2 (5 m-29)}{m^2-49} & y=-\frac{83-11 m}{m^2-49} \\ \end{array} \right)\\ m=\pm 7\implies \text {no solution} \]

OpenStudy (anonymous):

I got those answers, except I don't know how to put them into the form the question asks.

OpenStudy (anonymous):

\[ m=\pm 7\implies \text{no solutions}\\ m\ne \pm 7\implies \text{ unique solution}\\ m=\frac{47}4 \implies \text{ double solution } \]

OpenStudy (anonymous):

Where did you get the 47/4?